英語 (English)

発信型英語 (Communicative English) (歯 ((歯・口)A2 年))

マーシェソジェラルド · 准教授/ 鳴門教育大学 2 単位前期水 5 ・ 6

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】 The purpose of this class is to give students an opportunity to use the skills they have developed in the past, for practical and communicative purposes

【授業の概要】Students will complete a variety of tasks which follow several themes through the course of the term. Themes include ’talking about cities and places around the world’, ’vacations students have taken or would like to take’ and ’describing events and drawing conclusions from evidence’.

【キーワード】 [キーワード] 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】 The purpose of this class is to improve English communicative and

productive skills. Students will develop confidence as they use English for a variety of functions. Grammar and vocabulary will be dealt with sometimes, but they will arise in context, from students’ need to communicate as they complete tasks, rather than being introduced by the teacher without context. Language form will not be viewed as generative, but subservient to communication.

【授業の計画】 1. Introductions / Getting to know your fellow students and teacher

2. Pronunciation workshop

3. Using past tenses to describe your life

4. Describing other people’s life and abilities

5. Presentation preparation

6. Presentations

7. Using present tense to talk about University life

8. Comparing Universities in Canada and Japan

9. Using adjectives to talk about neighborhoods in Tokushima

10. Using adjectives to describe various cities

11. Describing your hometown

12. Describing cities and countries around the world

13. Describing neighborhoods in Toronto 14. Choosing an apartment in Toronto (group decision task)

15. Test

16. Final class 【教科書】 Students will use material prepared by the instructor and self-generated

material. There will be no textbook. 【参考書等】 [参考資料] 【成績評価の方法】 Speaking test: 25%, Presentation: 25%, Homework and participation:

30%, Attendance: 20% 【再試験の有無】 Students who miss the exam or fail to perform adequately will

not receive a reexamination except under special circumstances. Sufficient documentation must be provided.

【受講者へのメッセージ】 This is a practical English class, which means the focus will be on communication. You have already learned a lot of grammar but most of you have not had much opportunity to use what you have learned. In this class, you will be required to actually use the English which you have learned in the past. You have all got a lot of English ’packed’ inside your head. My job as your teacher is to help you get it out; to give you opportunities to speak. I will teach grammar and vocabulary sometimes, but not usually. Many grammar and vocabulary points will come up in the lesson, as you complete work and talk with your classmates. It is very important to ask questions in class, since this will raise important grammar and vocabulary points.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220896 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】 ⇒ マーシェソ .

English

Communicative English (歯 ((歯・口)A2 年))

Gerard Marchesseau · Associate Professor / Naruto University of Education 2 units 前期水 5 ・ 6

(平成 19 年度以前の授業科目: 『発信型英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ The purpose of this class is to give students an opportunity to use the

skills they have developed in the past, for practical and communicative purposes Outline⟩ Students will complete a variety of tasks which follow several themes

through the course of the term. Themes include ’talking about cities and places around the world’, ’vacations students have taken or would like to take’ and ’describing events and drawing conclusions from evidence’.

�������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ The purpose of this class is to improve English communicative and

productive skills. Students will develop confidence as they use English for a variety of functions. Grammar and vocabulary will be dealt with sometimes, but they will arise in context, from students’ need to communicate as they complete tasks, rather than being introduced by the teacher without context. Language form will not be viewed as generative, but subservient to communication.

Schedule⟩

1. Introductions / Getting to know your fellow students and teacher

2. Pronunciation workshop

3. Using past tenses to describe your life

4. Describing other people’s life and abilities

5. Presentation preparation

6. Presentations

7. Using present tense to talk about University life

8. Comparing Universities in Canada and Japan

9. Using adjectives to talk about neighborhoods in Tokushima

10. Using adjectives to describe various cities

11. Describing your hometown

12. Describing cities and countries around the world

13. Describing neighborhoods in Toronto

14. Choosing an apartment in Toronto (group decision task)

15. Test

16. Final class Textbook⟩ Students will use material prepared by the instructor and self-generated

material. There will be no textbook. Reference⟩ [参考資料] Evaluation Criteria⟩ Speaking test: 25%, Presentation: 25%, Homework and

participation:30%, Attendance: 20% Re-evaluation⟩ Students who miss the exam or fail to perform adequately will

not receive a reexamination except under special circumstances. Sufficient documentation must be provided.

Message⟩ This is a practical English class, which means the focus will be on communication. You have already learned a lot of grammar but most of you have not had much opportunity to use what you have learned. In this class, you will be required to actually use the English which you have learned in the past. You have all got a lot of English ’packed’ inside your head. My job as your teacher is to help you get it out; to give you opportunities to speak. I will teach grammar and vocabulary sometimes, but not usually. Many grammar and vocabulary points will come up in the lesson, as you complete work and talk with your classmates. It is very important to ask questions in class, since this will raise important grammar and vocabulary points.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220896 Contact (Office-Hour, Room, E-mail)⟩

⇒ Marchesseau .